# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 7 Numerical Verification Beyond Over-parameterized Matrix Sensing
In this section, we empirically examine the performance of our algorithm in over-parameterized models beyond the symmetric matrix sensing setting. Specifically, we focus on two problem classes: low-rank matrix recovery (in both symmetric and asymmetric forms) and sparse recovery. Our analysis centers on IPGD+, as we conjecture that all considered instances satisfy the $\mathcal{M}$-strict saddle property and $\mathcal{M}$-regularity conditions necessary for the improved convergence guarantees of IPGD+.
@@
The objective in low-rank matrix recovery is to recover a ground-truth rank-$r$ matrix ${\mathbf{M}^{\star}}$ under various observation models. One such example is the symmetric matrix sensing model (discussed in Section 6), where the observation operator satisfies the RIP. However, other examples of low-rank matrix recovery, such as matrix completion, may not satisfy RIP.
@@
In our experiments, we consider both symmetric and asymmetric variants of this problem. In the symmetric low-rank matrix recovery, the ground truth ${\mathbf{M}^{\star}}\in\mathbb{R}^{n\times n}$ is assumed to be symmetric and PSD. As mentioned in Section 1, it is common to model ${\mathbf{M}^{\star}}$ as ${\mathbf{M}^{\star}}=\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}$ and $r^{\prime}\geq r$, and directly optimize the lifted loss $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$. For this family of problems, we take the implicit region as =$\mathcal{M}=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}\mid\mathrm{rank}(\mathbf{X})\leq r\}$. We emphasize that this definition of $\mathcal{M}$ is similar to that provided in Equation 78, with the key difference that the constraints $\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ and $\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$ are omitted. This simpler formulation of the implicit region is more broadly applicable across various low-rank matrix recovery problems and, notably, does not depend on the specific observation model used (such as the RIP). Even within the aforementioned matrix sensing setting, we have observed in practice that the omitted constraints do not become active. In other words, $\mathbf{X}^{\sharp}_{t}=\mathcal{P}_{\mathcal{M}}(\mathbf{X}_{t})$ automatically satisfies $\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}^{\sharp}_{t}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ and $\left\lVert\mathbf{X}^{\sharp}_{t}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$ for almost all iterations $t$.
@@
In the asymmetric case, the target rank-$r$ matrix ${\mathbf{M}^{\star}}\in\mathbb{R}^{n_{1}\times n_{2}}$ is assumed to be potentially indefinite and rectangular. In this setting, it is common to employ a two-factor model ${\mathbf{M}^{\star}}=\mathbf{X}\mathbf{Y}^{\top}$, where $\mathbf{X}\in\mathbb{R}^{n_{1}\times r^{\prime}}$ and $\mathbf{Y}\in\mathbb{R}^{n_{2}\times r^{\prime}}$ with $r^{\prime}\geq r$, and, as in the symmetric case, to directly optimize the lifted loss $f(\mathbf{X},\mathbf{Y})=L(\mathbf{X}\mathbf{Y}^{\top})$. Correspondingly, the implicit region generalizes to =$\mathcal{M}=\{(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{n_{1}\times r^{\prime}}\times\mathbb{R}^{n_{2}\times r^{\prime}}\mid\mathrm{rank}(\mathbf{X})\leq r,\ \mathrm{rank}(\mathbf{Y})\leq r\}$. To promote balancedness in the obtained solution, it is also common to include an additional regularization term, $\|\mathbf{X}^{\top}\mathbf{X}-\mathbf{Y}^{\top}\mathbf{Y}\|_{\mathrm{F}}^{2}$ [42, 61].
In addition to the matrix sensing, we consider the following observation models:
  • Matrix completion [62, 63]: Here, the goal is to recover a rank-$r$ matrix ${\mathbf{M}^{\star}}$ based on a limited number of observed entries. In the symmetric variant, the recovery is performed by minimizing the objective function
    $f(\mathbf{X})=\sum_{(i,j)\in\Omega}\left((\mathbf{X}\mathbf{X}^{\top})_{i,j}-{\mathbf{M}^{\star}}_{i,j}\right)^{2},$ (123)
    where $\Omega\subseteq[n]\times[n]$ denotes the set of observed indices. Given an observation probability $10<p<$ 1, we generate $\Omega$ randomly as a uniformly selected subset of $[n]\times[n]$ with $|\Omega|=np$. In all of our experiments, we set $p=0.8$. For the asymmetric variant, we replace the symmetric factorization with the two-factor model and incorporate the balancing regularization term.
  • 1-bit matrix completion [64]: This is a generalization of the classical matrix completion problem, where we observe the entries of ${\mathbf{M}^{\star}}$ only through 1-bit (binary) measurements. Specifically, each measurement equals $1$ with probability $\sigma(\mathbf{M}^{\star}_{i,j})$ and 00 with probability $1-\sigma(\mathbf{M}^{\star}_{i,j})$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the logistic (sigmoid) function. Let $\alpha_{ij}$ denote the fraction of measurements for entry $(i,j)$ that are observed as $1$. The objective is to minimize
    $f(\mathbf{X})=\sum_{i,j=1}^{N}\left(\log\left(1+\exp\left(\left[\mathbf{X}\mathbf{X}^{\top}\right]_{i,j}\right)\right)-\alpha_{ij}\left[\mathbf{X}\mathbf{X}^{\top}\right]_{i,j}\right),$ (124)
    In our experiments, we set $\alpha_{ij}$ to its population expectation, i.e., $\alpha_{ij}=\sigma(\mathbf{M}^{\star}_{i,j})$. For the asymmetric case, we replace the symmetric factorization with the two-factor model and include the balancing regularization, following the same approach as in the previous task.
@@
In all of our experiments, we set $n_{1}=n_{2}=n=20$, $r^{\prime}=20$ (corresponding to a fully over-parameterized setting), and $r=3$. For the symmetric variants, the ground-truth matrix is constructed as $\mathbf{M}^{\star}=\mathbf{U}^{\star}\mathbf{\Sigma}^{\star}\mathbf{U}^{\star\top}$, where $\mathbf{U}^{\star}\in\mathbb{R}^{n\times r}$ is a randomly generated orthogonal matrix, and $\mathbf{\Sigma}^{\star}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with diagonal entries $10$, $5$, and $1$. Similarly, for the asymmetric variants, the ground-truth matrix is constructed as $\mathbf{M}^{\star}=\mathbf{U}^{\star}\mathbf{\Sigma}^{\star}\mathbf{V}^{\star\top}$, where $\mathbf{U}^{\star}\in\mathbb{R}^{n_{1}\times r}$ and $\mathbf{V}^{\star}\in\mathbb{R}^{n_{2}\times r}$ are randomly generated orthogonal matrices, and $\mathbf{\Sigma}^{\star}$ is the same diagonal matrix as in the symmetric case. In all experiments, the initial points are initialized as all-zero matrices. The perturbation radius is set to $\gamma=1\times 10^{-15}$, and the gradient norm threshold is set to $G=1\times 10^{-7}$. We use $F=1\times 10^{-10}$ and $T_{\mathrm{escape}}=50$ in all settings. The step size is set to $\eta=10^{-1}$ for matrix sensing and 1-bit matrix completion, and to $\eta=2\times 10^{-2}$ for matrix completion.
@@
Figure 4 illustrates the behavior of IPGD+ (Algorithm 2) applied to the symmetric and asymmetric variants of matrix sensing, matrix completion, and 1-bit matrix completion. Our first key observation is that IPGD+ not only stays close to the defined implicit region $\mathcal{M}$—as evidenced by the consistently small residual norm—but also efficiently converges to a ground-truth solution in $\mathcal{X}_{\mathcal{M}}$. Notably, the final distance to $\mathcal{X}_{\mathcal{M}}$ precisely matches the residual norm across all instances considered. This phenomenon indicates that our signal-residual decomposition provides an accurate characterization of the IPGD+ trajectory. Moreover, a closer examination of Figure 4 reveals that the deviation rate $r(\cdot)$ (Definition 5) ultimately governs the final residual norm. Specifically, it suggests that a larger cumulative deviation rate, quantified as $\bar{r}_{T}=\sum_{t=0}^{T-1}r(\mathbf{x}_{t})$, leads to a faster growth in the residual norm. This observation aligns with the bound established in Equation 13. Later in this section, we provide a more detailed analysis of the relationship between the cumulative deviation rate and the final residual norm.
@@
Sparse recovery
In this problem, the goal is to recover a $s$-sparse vector $\bm{\theta}^{\star}\in\mathbb{R}^{d}$ from a dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}=\left\langle\bm{\theta}^{\star},\mathbf{x}_{i}\right\rangle$. We model the ground-truth vector as a two-layer decomposition, $\bm{\theta}^{\star}=\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v}$, and minimize the objective function
$f(\mathbf{u},\mathbf{v})=\sum_{i=1}^{N}\left(y_{i}-\left\langle\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v},\mathbf{x}_{i}\right\rangle\right)^{2}.$ (125)
Notably, under the decomposition $\bm{\theta}^{\star}=\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v}$, the positive entries of $\bm{\theta}^{\star}$ are represented solely by the first layer $\mathbf{u}\odot\mathbf{u}$, while the negative entries are captured only by the second layer $-\mathbf{v}\odot\mathbf{v}$. Based on this observation, we define the implicit region as =$\mathcal{M}=\left\{(\mathbf{u},\mathbf{v})\mid\mathbf{u}_{i}=0\;\forall i\notin\mathcal{S}_{+},\;\mathbf{v}_{i}=0\;\forall i\notin\mathcal{S}_{-}\right\},$ where $\mathcal{S}_{+}=\{i\mid\bm{\theta}^{\star}_{i}>0\}$ and $\mathcal{S}_{-}=\{i\mid\bm{\theta}^{\star}_{i}<0\}$ collect the indices of the positive and negative entries of $\bm{\theta}^{\star}$.
In our experiments, we set $d=150$, $N=300$, and $k=5$. The ground-truth vector is set to an all-zero vector, except for its first five entries, which are set to $10$, $-5$, $3$, $-2$, and $1$, respectively. The entries of $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are independently drawn from a standard normal distribution. As in previous case studies, the initialization is set to an all-zero vector, with a perturbation radius of $\gamma=1\times 10^{-15}$. We set the step size to $\eta=1\times 10^{-2}$, the gradient norm threshold to $G=1\times 10^{-7}$, the escape time to $T_{\mathrm{escape}}=50$ and the termination threshold to $F=1\times 10^{-10}$.
@@
8 Conclusion and Future Directions
In this paper, we study the question of why gradient-based methods in over-parameterized models often yield low-dimensional solutions, even without explicit constraints or regularization. To provide a partial answer to this long-standing question, we study the conditions that enable implicit regularization by investigating when gradient-based methods converge to second-order stationary points (SOSPs) within an implicit low-dimensional region of a smooth and nonconvex function. We identify and formalize the key conditions under which such implicit regularization arises: namely, the combined ability to escape strict saddle points using infinitesimal perturbations and to remain close to an implicit low-dimensional region by controlling the deviation rate. Building on these insights, we introduce infinitesimally perturbed gradient descent (IPGD) algorithm and show that it can provably satisfy both conditions. We further study the application of IPGD in the context of over-parameterized matrix sensing. Our theoretical findings are further supported by extensive empirical results, suggesting that these principles extend more broadly across diverse over-parameterized learning settings.
We now discuss two future directions that we believe our proposed framework could potentially address.
@@ 8 Conclusion and Future Directions
While this work focuses on GD with infinitesimal perturbations, a natural follow-up question is whether our framework extends to other widely used variants of GD, such as preconditioned [65, 66, 67, 68], accelerated [28, 29], or stochastic gradient descent (SGD) [21]. Take SGD, for instance: unlike IPGD, the perturbations in SGD need not be small; however, they are also not isotropic, meaning they affect the signal and residual norms unevenly. In some cases, such anisotropic perturbations may be advantageous; for example, when they impact the residual norm much less than the signal norm. An intriguing open question is whether the proposed framework, based on the signal-residual decomposition and the deviation rate from the implicit region, can be generalized to encompass such settings. Developing a broader notion of deviation rate could enable direct, principled comparisons between algorithms.
@@
One notable limitation of our current framework is its reliance on the gradient and Hessian Lipschitz properties of the objective function within a neighborhood of the implicit region, restricting its applicability to smooth problems. An enticing direction for future work is to explore whether our framework can be extended to nonsmooth settings by leveraging tools from variational analysis. Such an extension would significantly broaden the scope of our framework, encompassing important applications such as robust principal component analysis, robust matrix recovery, and neural networks with nonsmooth activation functions. We emphasize, however, that such an extension is far from trivial, as the challenge of escaping strict saddle points is substantially more delicate in nonsmooth settings. Notably, recent work has shown that even in exactly-parameterized problems, first-order methods such as the stochastic sub-gradient method [69] and proximal methods [18] can only escape active strict saddle points. This distinction introduces an additional layer of complexity for developing a theoretical understanding of implicit regularization in nonsmooth regimes.

### v2

@@ 7 Numerical Verification Beyond Over-parameterized Matrix Sensing
In this section, we empirically examine the performance of our algorithm in over-parameterized models beyond the symmetric matrix sensing setting. Specifically, we focus on two problem classes: low-rank matrix recovery (in both symmetric and asymmetric forms) and sparse recovery. Throughout all experiments, we employ IPGD+, which consistently demonstrates a linear convergence rate in practice.
@@
Recall that the goal in low-rank matrix recovery is to recover a ground-truth rank-$r$ matrix $\mathbf{\Theta}^{\star}$ under various loss functions. We consider both symmetric and asymmetric variants. In the symmetric low-rank matrix recovery, the ground truth matrix $\mathbf{\Theta}^{\star}\in\mathbb{R}^{n\times n}$ is assumed to be PSD. We consider the reparameterization map $\mathbf{\Theta}=\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}$ and $r^{\prime}\geq r$, and optimize the loss function $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$. For this family of problems, we take the implicit region as $\mathcal{M}=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r\}$. We emphasize that this definition of $\mathcal{M}$ is similar to Equation 79, with the key difference that the constraints $\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ and $\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$ are omitted. This simpler formulation of the implicit region is more broadly applicable across various low-rank matrix recovery problems and, notably, does not depend on the specific observation model used (such as the RIP). Even within the aforementioned matrix sensing setting, we have observed in practice that the omitted constraints do not become active. In other words, $\mathbf{X}^{\sharp}_{t}=\mathcal{P}_{\mathcal{M}}(\mathbf{X}_{t})$ automatically satisfies $\big\|\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}^{\sharp}_{t}}\big\|\leq 4\sqrt{r}\kappa\delta$ and $\big\|\mathbf{X}^{\sharp}_{t}\big\|\leq 2\sqrt{\sigma_{1}^{\star}}$ for almost all iterations $t$.
@@
In the asymmetric case, the target rank-$r$ matrix $\mathbf{\Theta}^{\star}\in\mathbb{R}^{n_{1}\times n_{2}}$ is assumed to be potentially indefinite and rectangular. We consider the reparameterization map $\mathbf{\Theta}=\mathbf{X}\mathbf{Y}^{\top}$, where $\mathbf{X}\in\mathbb{R}^{n_{1}\times r^{\prime}}$ and $\mathbf{Y}\in\mathbb{R}^{n_{2}\times r^{\prime}}$ with $r^{\prime}\geq r$, and optimize the loss function $f(\mathbf{X},\mathbf{Y})=L(\mathbf{X}\mathbf{Y}^{\top})$. We take the rank-balanced implicit region as $\mathcal{M}=\{(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{n_{1}\times r^{\prime}}\times\mathbb{R}^{n_{2}\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r,\ \mathrm{rank}(\mathbf{Y})\leq r\}$.
Beyond matrix sensing, we also consider the following two settings:
  • Matrix completion [johnson1990matrix, candes2012exact]: Our goal is to recover a rank-$r$ matrix $\mathbf{\Theta}^{\star}$ from a subset of its observed entries $\Omega\subseteq[n]\times[n]$. The loss function is defined as
    $L(\mathbf{\Theta})=\sum_{(i,j)\in\Omega}\left(\mathbf{\Theta}_{i,j}-\mathbf{\Theta}^{\star}_{i,j}\right)^{2}.$ (124)
    We generate $\Omega$ uniformly at random with $|\Omega|=np$, where the observation probability is fixed at $p=0.8$ in all experiments.
  • 1-bit matrix completion [davenport20141]: This problem extends matrix completion by observing only binary (1-bit) measurements of the underlying matrix $\mathbf{\Theta}^{\star}$. Each entry $(i,j)$ is observed as $1$ with probability $\sigma(\mathbf{\Theta}^{\star}_{i,j})$ and 0 otherwise, Here, $\sigma(x)=(1+e^{-x})^{-1}$ denotes the sigmoid function. The corresponding expected loss function is
    $L(\mathbf{\Theta})=\sum_{i,j}\left(\log\left(1+\exp\left(\mathbf{\Theta}_{i,j}\right)\right)-\sigma(\mathbf{\Theta}^{\star}_{i,j})\mathbf{\Theta}_{i,j}\right).$ (125)
@@
In all of our experiments, we set $n_{1}=n_{2}=n=20$, $r^{\prime}=20$, and $r=3$. For the symmetric variants, the ground-truth matrix is constructed as $\mathbf{\Theta}^{\star}=\mathbf{U}^{\star}\mathbf{\Sigma}^{\star}\mathbf{U}^{\star\top}$, where $\mathbf{U}^{\star}\in\mathbb{R}^{n\times r}$ is a randomly generated orthogonal matrix, and $\mathbf{\Sigma}^{\star}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with entries $10$, 55, and $1$. Similarly, for the asymmetric variants, the ground-truth matrix is constructed as $\mathbf{\Theta}^{\star}=\mathbf{U}^{\star}\mathbf{\Sigma}^{\star}\mathbf{V}^{\star\top}$, where $\mathbf{U}^{\star}\in\mathbb{R}^{n_{1}\times r}$ and $\mathbf{V}^{\star}\in\mathbb{R}^{n_{2}\times r}$ are randomly generated orthogonal matrices, and $\mathbf{\Sigma}^{\star}$ is the same diagonal matrix as in the symmetric case. In all experiments, the initial points are initialized as all-zero matrices. The perturbation radius is set to $\gamma=1\times 10^{-15}$, and the gradient norm threshold is set to $G=1\times 10^{-7}$. We use $F=1\times 10^{-10}$ and $T_{\mathrm{escape}}=50$ in all settings. The step size is set to $\eta=0.1$ for matrix sensing and 1-bit matrix completion, and to $\eta=0.02$ for matrix completion.
@@
Figure 4 presents the behavior of IPGD+ (Algorithm 2) applied to the symmetric and asymmetric variants of matrix sensing, matrix completion, and 1-bit matrix completion. Our first key observation is that IPGD+ not only stays close to the defined implicit region $\mathcal{M}$—as evidenced by the consistently small residual norm—but also efficiently converges to a ground-truth solution in $\mathcal{X}_{\mathcal{M}}$. Notably, the final distance to $\mathcal{X}_{\mathcal{M}}$ precisely matches the residual norm across all instances considered. This phenomenon indicates that our signal-residual decomposition provides an accurate characterization of the performance of IPGD+. Moreover, a closer examination of Figure 4 reveals that the cumulative effect of the deviation rate $r(\cdot)$ (Definition 5) ultimately governs the final residual norm. Specifically, it suggests that a larger cumulative deviation rate, quantified as $\bar{r}_{T}=\sum_{t=0}^{T-1}r(\mathbf{x}_{t})$, leads to a faster growth in the residual norm. This observation aligns with the bound established in Equation 14. Later in this section, we provide a more detailed analysis of the relationship between the cumulative deviation rate and the final residual norm.
@@
Sparse recovery
In this problem, the goal is to recover a $k$-sparse vector $\bm{\theta}^{\star}\in\mathbb{R}^{d}$ from a dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}=\left\langle\bm{\theta}^{\star},\mathbf{x}_{i}\right\rangle$. We use the over-parameterized model $\bm{\theta}^{\star}=\mathbf{u}\odot\mathbf{v}$, and minimize the objective function
$f(\mathbf{u},\mathbf{v})=\sum_{i=1}^{N}\left(y_{i}-\left\langle\mathbf{u}\odot\mathbf{v},\mathbf{x}_{i}\right\rangle\right)^{2}.$ (126)
We use the implicit region defined in Example 2: $\mathcal{M}_{\text{balanced}}=\{(\mathbf{u},\mathbf{v}):\mathbf{u}_{i}=\mathbf{v}_{i}=0\ \text{if}\ \bm{\theta}^{\star}_{i}=0\}$.
@@
In our experiments, we set $d=150$, $N=300$, and $k=5$. The ground-truth vector is set to an all-zero vector, except for its first five entries, which are set to $10$, 5-5, 33, 2-2, and $1$, respectively. The entries of $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are independently drawn from a standard normal distribution. As in previous case studies, the initialization is set to an all-zero vector, with a perturbation radius of $\gamma=1\times 10^{-15}$. We set the step size to $\eta=1\times 10^{-2}$, the gradient norm threshold to $G=1\times 10^{-7}$, the escape time to $T_{\mathrm{escape}}=50$ and the termination threshold to $F=1\times 10^{-10}$.
@@
8 Conclusion
In this paper, we study why gradient-based methods in over-parameterized models often enjoy implicit regularization towards low-dimensional solutions, even in the absence of explicit constraints or regularization. To address this question, we investigate the conditions that enable implicit regularization by examining when gradient descent converges to a second-order stationary point (SOSP) within an implicit low-dimensional region of a smooth, nonconvex function. We identify and formalize the key conditions under which such implicit regularization emerges: namely, the combined ability to (i) initialize within a neighborhood of the implicit region, (ii) escape strict saddle points through infinitesimal perturbations, and (iii) remain close to the implicit region by controlling the deviation rate. Building on these insights, we introduce the infinitesimally perturbed gradient descent (IPGD) algorithm and show that, with proper initialization and reparameterization, it can provably satisfy all of these conditions. We further analyze the application of IPGD in the context of over-parameterized matrix sensing. Our theoretical results are supported by extensive empirical evidence, suggesting that these principles extend broadly across diverse over-parameterized learning settings.
We now discuss two future directions that our proposed framework could potentially address.
@@ 8 Conclusion
While this work focuses on GD with infinitesimal perturbations, a natural follow-up question is whether our framework extends to other widely used variants of GD, such as preconditioned [zhang2021preconditioned, zhang2023preconditioned, tong2021accelerating, xu2023power], accelerated [carmon2018accelerated, agarwal2017finding], or stochastic gradient descent (SGD) [fang2019sharp]. Take SGD, for instance: unlike IPGD, the perturbations in SGD need not be small; however, they are also not isotropic, meaning they affect the signal and residual norms unevenly. In some cases, such anisotropic perturbations may be advantageous; for example, when they impact the residual norm much less than the signal norm. An intriguing open question is whether the proposed framework, based on the signal-residual decomposition and the deviation rate from the implicit region, can be generalized to encompass such settings. Developing a broader notion of deviation rate could enable direct, principled comparisons between algorithms.
@@
One notable limitation of our current framework is its reliance on the gradient and Hessian Lipschitz properties of the objective function within a neighborhood of the implicit region, restricting its applicability to smooth problems. An enticing direction for future work is to explore whether our framework can be extended to nonsmooth settings by leveraging tools from variational analysis. Such an extension would significantly broaden the scope of our framework, encompassing important applications such as robust principal component analysis, robust matrix recovery, and neural networks with nonsmooth activation functions. We emphasize, however, that such an extension is far from trivial, as the challenge of escaping strict saddle points is substantially more delicate in nonsmooth settings. Notably, recent work has shown that even in exactly-parameterized problems, first-order methods such as the stochastic sub-gradient method [davis2025active] and proximal methods [davis2019active] can only escape active strict saddle points. This distinction introduces an additional layer of complexity for developing a theoretical understanding of implicit regularization in nonsmooth regimes.